OpenStudy (anonymous):

At time t>=0, the velocity of a particle moving along a straight line is v(t)=t^2-4t+3. a) Find the particle's acceleration each time the velocity is zero. b) When is the particle moving forward? c) When is the particle's velocity decreasing

OpenStudy (anonymous):

wish i could read, that is the velocity. ok interpret the first question as set \[t^2-4t+3=0\] solve for \(t\)

OpenStudy (anonymous):

easy for this one since it factors as \[(t-1)(t-3)=0\]

OpenStudy (anonymous):

2) moving forward means \(v(t)>0\) and since this is a parabola that opens up, it is negative between the roots and positive outside them

OpenStudy (anonymous):

3) take the derivative, get a line. solve for where the line is negative, aka below the \(x\) axis

OpenStudy (anonymous):

So at t=1, t=3 is the acceleration?

OpenStudy (anonymous):

no, that is where the velocity is zero that is the answer to question 1

OpenStudy (anonymous):

oh man i really can't read jesus, ignore everything i wrote it is asking about ACCELERATION ! sorry, let me delete everything, put on a dunce cap and start again

OpenStudy (anonymous):

haha its okay

OpenStudy (anonymous):

velocity is \(v(t)=t^2-4t+3\) and it is zero at \(t=1,t=3\) at least we got that right

OpenStudy (anonymous):

would I just plug in V(1) and v(3)?

OpenStudy (anonymous):

acceleration is \(a(t)=2t-4\) and the first question asks what is acceleration when the velocity is \(0\) to take \(a(1)\) and \(a(3)\)

OpenStudy (anonymous):

plug in to the derivative for acceleration if you plug back in to \(v\) you get zero, because that is how you solved for \(1\) and \(3\) to begin with

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

now for question two my answer was actually correct, because it asks when it is moving forward and backward. \(v(t)=t^2-4t+3\) is a parabola facing up so it is negative between \(1\) and \(3\) and positive outside , i.e. when \(t<1\) and when \(t>3\)

OpenStudy (anonymous):

for question 3, velocity is decreasing when acceleration is negative, so solve \[2t-4<0\] which should be easy enough sorry i messed up so bad, my reading comprehension is not what it should be

OpenStudy (anonymous):

its okay. Thanks a lot! Im going to solve it out and let you know what I get